OpenStudy (anonymous):

Find the first four terms in the binomial expansion of (2+x)^0.5

OpenStudy (zpupster):

what if it were 2+x)^2

OpenStudy (anonymous):

well then you could just square the bracket, you wouldn't need to use binomial at all.

OpenStudy (anonymous):

(2+x)^2 = 4 + 4x + x^2

OpenStudy (zpupster):

sqrt(x+2)

OpenStudy (anonymous):

Ok, so how do you do that using binomial theorem please?

OpenStudy (zpupster):

\[(1+x)^{a}=1+ax+a(a-1)\,\frac{x^{2}}{2!}+a(a-1)(a-2)\,\frac{x^{3}}{3!}\]

OpenStudy (zpupster):

that should give you your first 4

OpenStudy (anonymous):

Thank you